# A Survey of AIOps for Failure Management in the Era of Large Language Models

### v3

@@ 1.1. Why are LLMs Beneficial for AIOps on Failure Management?
  • Need for complex feature extraction engineering. These AIOps methods typically require extensive preprocessing and feature engineering to extract useful information from raw data. They have limited capabilities in understanding and processing data, especially in handling unstructured data such as logs and traces, which appears relatively weak.
@@ 2.1. Data Source for AIOps
  • Logs. Logs are detailed records of events that occur within the system. They can include error messages, transaction records, user activity, and system operations. Logs are essential for diagnosing issues, understanding system behavior, and tracking changes over time.
@@ 2.2. LLM-based Approaches for AIOps
  • Parameter-Efficient Fine-Tuning. This includes techniques such as Layer-Freezing, Adapter Tuning, and Task-Conditional Fine-Tuning. These methods require tuning only a small subset of the model’s parameters, making them more efficient and practical for AIOps applications. Layer-Freezing involves freezing most of the model’s layers and only fine-tuning the top layers, allowing for significant reductions in computational cost and training time. Adapter Tuning introduces small, trainable adapter modules within each layer of the pre-trained model, enabling the model to adapt to new tasks with minimal. Task-Conditional Fine-Tuning involves adding task-specific output heads to the model’s output and training them accordingly. Parameter-efficient fine-tuning is the predominant approach in the AIOps field.
@@ 4.1. Failure Prediction
Formally, given a time series of historical system data $(X=\{x_{1},x_{2},\ldots,x_{t}\}$, the goal is to predict the probability $P(f_{t+\Delta t})$ of a failure $f$ occurring within a future interval $\Delta t$. This process can be expressed as $P(f_{t+\Delta t}|X)$, which is the predicted probability of a failure occurring at time $t+\Delta t$ given the historical data $X$.
@@ 4.2. Anomaly Detection
Assume the actual observed values $X_{actual}=\{x_{T+1},x_{T+2},...,x_{T+\tau}\}$. The predicted values $\hat{X}$ are then compared with it. An anomaly is flagged if the deviation $\delta$ between $\hat{X}$ and $X_{actual}$ exceeds a predefined threshold $\epsilon$.
@@ 5.1. Failure Localization
In the era of large language models (LLMs), this type of work is relatively rare, and the identified failure types to be located are more diverse. Komal et al.(Sarda et al., 2023; Sarda, 2023) leverage zero-shot prompt engineering approaches based on metrics data to identify problematic host names, pod names, and deployment names. Devjeet et al.(Roy et al., 2024) utilize incident reports with React (Yao et al., 2022) for root cause analysis in an out-of-domain setting on a static dataset of real-world production incidents, ultimately identifying the problematic task names. RealTCD (Li et al., 2024b) leverages domain knowledge to discover temporal causal relationships for root cause analysis without interventional targets, introducing LLM-guided meta-initialization to extract meta-knowledge from textual information hidden in systems to enhance discovery quality. LogConfigLocalizer (Shan et al., 2024) proposes an LLM-based two-stage strategy for end-users to localize root-cause configuration properties based on logs. mABC(Zhang et al., 2024b) utilizes alert data (a processed form of metrics data) and employs a multi-agent blockchain-inspired collaboration to identify the specific node where the anomaly occurred.
@@ 6.4. Script Generation
The aforementioned command recommendation methods still require OCEs to write some commands or input natural language instructions, which involves a certain level of human effort. Therefore, many works have begun to directly research the generation of complete remediation scripts for OCEs to execute.

### v4

@@ 1.1. Why are LLMs Beneficial for AIOps on Failure Management?
  • Need for complex feature extraction engineering. These AIOps methods typically require extensive preprocessing and feature engineering to extract useful information from raw data. They have limited capabilities in understanding and processing data, especially in handling unstructured data such as logs and traces, which appearing relatively weak.
@@ 2.1. Data Source for AIOps
  • Logs. Logs are detailed records of events that occur within the system. They can include error messages, transaction records, user activities, and system operations. Logs are essential for diagnosing issues, understanding system behavior, and tracking changes over time.
@@ 2.2. LLM-based Approaches for AIOps
  • Parameter-Efficient Fine-Tuning. This includes techniques such as Layer-Freezing, Adapter Tuning, and Task-Conditional Fine-Tuning. These methods require tuning only a small subset of the model’s parameters, making them more efficient and practical for AIOps applications. Layer-Freezing involves freezing most of the model’s layers and only fine-tuning the top layers, allowing for significant reductions in computational cost and training time. Adapter Tuning introduces small, trainable adapter modules within each layer of the pre-trained model, enabling the model to adapt to new tasks with minimal effort. Task-Conditional Fine-Tuning involves adding task-specific output heads to the model’s output and training them accordingly. Parameter-efficient fine-tuning is the predominant approach in the AIOps field.
@@ 4.1. Failure Prediction
Formally, given a time series of historical system data $X=\{x_{1},x_{2},\ldots,x_{t}\}$, the goal is to predict the probability $P(f_{t+\Delta t})$ of a failure $f$ occurring within a future interval $\Delta t$. This process can be expressed as $P(f_{t+\Delta t}|X)$, which is the predicted probability of a failure occurring at time $t+\Delta t$ given the historical data $X$.
@@ 4.2. Anomaly Detection
Assume the actual observed values $X_{actual}=\{x_{T+1},x_{T+2},...,x_{T+\tau}\}$. The predicted values $\hat{X}$ are then compared with $X_{actual}$. An anomaly is flagged if the deviation $\delta$ between $\hat{X}$ and $X_{actual}$ exceeds a predefined threshold $\epsilon$.
@@ 5.1. Failure Localization
In the era of large language models (LLMs), this type of work is relatively rare, and the identified failure types to be located are more diverse. Komal et al. (Sarda et al., 2023; Sarda, 2023) leverage zero-shot prompt engineering approaches based on metrics data to identify problematic host names, and deployment names. Devjeet et al. (Roy et al., 2024) utilize incident reports with React (Yao et al., 2022) for root cause analysis in an out-of-domain setting on a static dataset of real-world production incidents, ultimately identifying the problematic task names. RealTCD (Li et al., 2024b) leverages domain knowledge to discover temporal causal relationships for root cause analysis without interventional targets, introducing LLM-guided meta-initialization to extract meta-knowledge from textual information hidden in systems to enhance discovery quality. LogConfigLocalizer (Shan et al., 2024) proposes an LLM-based two-stage strategy for end-users to localize root-cause configuration properties based on logs. mABC (Zhang et al., 2024b) utilizes alert data (a processed form of metrics data) and employs a multi-agent blockchain-inspired collaboration to identify the specific node where the anomaly occurred.
@@ 6.4. Script Generation
The aforementioned command recommendation methods still require OCEs to write some commands or input natural language instructions, which involving a certain level of human effort. Therefore, many works have begun to directly research the generation of complete remediation scripts for OCEs to execute.